# Geometric families of multiple elliptic Gamma functions and arithmetic applications, I

### v1

@@ Abstract:
This is the first paper in a series where we study arithmetic applications of the multiple elliptic Gamma functions originated from theoretical physics. The main purpose of this paper is the introduction of a framework for applications of these functions to Hilbert’s 12th problem for general number fields with exactly one complex place following recent work by Bergeron, Charollois and García. Namely, we define geometric families of the multiple elliptic Gamma functions, upgrading the construction carried out by Felder, Henriques, Rossi and Zhu for rank $3$ lattices to lattices of higher ranks. These functions enjoy transformation properties under an action of the special linear group $\mathrm{SL}_{n}(\mathbb{Z})$ for $n\geq 2$ involving some Bernoulli rational functions as their so-called modularity defect. A second purpose of this paper is to use this collection of Bernoulli rational functions to construct (n1)(n-1)-cocycles for specific subgroups of $\mathrm{SL}_{n}(\mathbb{Z})$ associated to units groups in totally real number fields and use these cocycles to compute partial zeta values at $s=0$.
@@ 1 Introduction
where as usual the notation $\widehat{\tau_{j}}$ indicates that the variable $\tau_{j}$ should be omited. In [FHRZ08] Felder, Henriques, Rossi and Zhu enriched the theory of the elliptic Gamma function by introducing geometric families of elliptic Gamma functions associated to arbitrary rank $3$ lattices, offering a comprehensive perspective of the underlying geometric phenomena. Namely, if $L$ is a rank $3$ lattice and $a,b\in\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$ are two primitive linear forms on $L$, they define general functions $\Gamma_{a,b}:\mathbb{C}\times\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{C})\to\mathbb{C}$ using Ruijsenaars’ elliptic $\Gamma$ function as a building block and show that the collection of functions obtained when varying $a,b$ enjoys nice properties under an action of $\mathrm{SL}_{3}(\mathbb{Z})$ on $L$. The two most important properties in that regard are the so-called modular and equivariance properties. For any linearly independent $a,b,c\in\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$, there is a rational function $P_{a,b,c}\in\mathbb{Q}[w](x)$ such that the equality:
@@
In the first part of this article, we upgrade the construction carried out in [FHRZ08] to higher degree $G_{r}$ functions (see section 2). For a lattice $\Lambda=\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$ of rank $n=r+2\geq 2$ we introduce a collection of geometric functions $G_{n-2,a_{1},\dots,a_{n-1}}$ attached to families of n1n-1 linear forms $a_{1},\dots,a_{n-1}\in\Lambda$ (see Proposition 2.2) which are built using Nishizawa’s $G_{n-2}$ functions. For rank $n=2,3$ lattices, we recover the functions $\theta_{a}=G_{0,a}(0)$ and $\Gamma_{a,b}=G_{1,a,b}(0)$ respectively. In section 2 we show that the collection of $G_{n-2,a_{1},\dots,a_{n-1}}$ functions obtained when varying the linear forms $a_{1},\dots,a_{n-1}$ satisfy modular and equivariance properties similar to those satisfied by the collections of $\theta_{a}$ and $\Gamma_{a,b}$ functions. Namely, our first main result is a general version of formulae (1) and (2) for the $G_{n-2,a_{1},\dots,a_{n-1}}$ functions involving a collection of higher degree Bernoulli rational functions $B_{n,a_{1},\dots,a_{n}}$ (see Definition 2.2).
@@
In applications the linear forms $a_{j}$ will be taken of the form $g_{j}\cdot a$ for some base point $a\in V^{\vee}$ and the invertible matrix $g_{j}$ in some subgroup $U$ of $\mathrm{SL}_{n}(\mathbb{Z})$. At the end of section 3 we give examples of specific subgroups $U$ of $\mathrm{SL}_{n}(\mathbb{Z})$ for which any family $g_{0},\dots,g_{n}\in U$ is such that the family $g_{0}\cdot a,\dots,g_{n}\cdot a$ satisfies the hypothesis of Corollary 1 for any base point $a\in V^{\vee}$. The collection of functions $(\phi_{n,a})_{a}$ thus reduces to a collection of homogeneous (n1)(n-1)-cocycles for $U$. In the arithmetic applications we have in mind to the computation of partial zeta values in totally real number fields at $s=0$, the group $U$ will arise from the group of totally positive units of a given totally real number field. Following [Shi76], [Col88], [DyDF14] and [CDG15] we will prove that the partial zeta values in totally real number fields may be expressed as combinations of values of the geometric Bernoulli functions $B_{n,a_{1},\dots,a_{n}}$ which appear in the study of the collection of $G_{n-2,a_{1},\dots,a_{n-1}}$ functions:
@@ 2 Geometric families of multiple elliptic Gamma functions
In this section, we recall the properties of the multiple elliptic Gamma functions (the $G_{r}$ functions) and construct geometric families $G_{r,a_{1},\dots,a_{r+1}}$ of these functions attached to a family of $r+1$ linear forms $a_{1},\dots,a_{r+1}$, upgrading the $G_{r}$ functions to collections of equivariant functions for $\mathrm{SL}_{r+2}(\mathbb{Z})$ by adapting the construction of the $\Gamma_{a,b}$ functions in [FHRZ08] to higher degrees.
@@ 2.1.3 Distribution relations
$\mathbb{Z}_{>0}-\{0,1/N,\dots,(N-1)/N\}=\mathbb{Z}_{>0}/N,\penalty 10000\ \penalty 10000\ \mathbb{Z}_{\geq 0}+\{0,1/N,\dots,(N-1)/N\}=\mathbb{Z}_{\geq 0}/N$
@@ 2.2 Geometric $G_{r,\underline{a}}$ functions
In this section we upgrade the construction carried out by Felder, Henriques, Rossi and Zhu to higher degree and define geometric families $G_{r,a_{1},\dots,a_{r+1}}$ parametrised by $r+1$ linear forms of the $G_{r}$ functions that encompass the geometric families $\Gamma_{a,b}$ of the elliptic Gamma function. We then prove Theorem 1, i.e. the modular property for the $G_{r,a_{1},\dots,a_{r+1}}$ functions and their equivariance property under the action of $\mathrm{SL}_{r+2}(\mathbb{Z})$. The construction of our $G_{r,a_{1},\dots,a_{r+1}}$ functions is adapted from the construction of the elliptic $\Gamma_{a,b}$ functions as we use a generalised version of the alternative definition given by [[FHRZ08], Proposition 3.5] and then revert the computations to prove that the definition is indeed valid.
@@
We may now define the geometric variants of the $G_{r}$ functions using lemma 2.2 and adapting [[FHRZ08], Proposition 3.5] to higher rank lattices:
@@
$G_{r,\underline{a}}^{\underline{\alpha}}(v)(w,x)^{(-1)^{D}}=\prod_{\delta\in F(\underline{a},\underline{\alpha},v)/\mathbb{Z}\gamma}\prod_{\underline{m}\,\geq 0}\Big[\left(1-e^{2i\pi\left(\sum_{j=1}^{r+1}\frac{(d_{j}m_{j}+(1+d_{j})/2)x(\alpha_{j})}{x(\gamma)}-\frac{w+x(\delta)}{x(\gamma)}\right)}\right)\\ \times\left(1-e^{2i\pi\left(\sum_{j=1}^{r+1}\frac{(d_{j}m_{j}+(1-d_{j})/2)x(\alpha_{j})}{x(\gamma)}+\frac{w+x(\delta)}{x(\gamma)}\right)}\right)^{(-1)^{r}}\Big]$
Let us denote $C^{+}(\underline{a},\underline{\alpha},v,x)$ for the set of $\delta^{\prime}\in v+L$ satisfying for all $1\leq j\leq r+1$:
@@
and similarly denote $C^{-}(\underline{a},\underline{\alpha},v,x)$ for the set of $\delta^{\prime}\in v+L$ satisfying for all $1\leq j\leq r+1$:
@@
Consider $\delta^{\prime}\in C^{+}(\underline{a},\underline{\alpha},v,x)$. This means that if $d_{j}=1$ then $a_{j}(\delta^{\prime})>0$ and we can perform Euclidian division by $a_{j}(\alpha_{j})$ so that there exists a unique integer $m_{j}>0$ such that $0\leq a_{j}(m_{j}\alpha_{j}-\delta^{\prime})<a_{j}(\alpha_{j})$. On the contrary, if $d_{j}=-1$ then $a_{j}(\delta^{\prime})\leq 0$ and there exists a unique integer $m_{j}\geq 0$ such that $0\leq a_{j}(-m_{j}\alpha_{j}-\delta^{\prime})<a_{j}(\alpha_{j})$. Then for all $1\leq k\leq r+1$:
$0\leq a_{k}\left(\sum_{j=1}^{r+1}d_{j}m_{j}\alpha_{j}-\delta^{\prime}\right)<a_{k}(\alpha_{k})$
This shows that $C^{+}(\underline{a},\underline{\alpha},v,x)$ can be written as a disjoint union:
$C^{+}(\underline{a},\underline{\alpha},v,x)=\bigcup_{\delta\in F(\underline{a},\underline{\alpha},v)}\bigcup_{\underline{m}\,\geq 0}\left\{-\delta+\sum_{j=1}^{r+1}(d_{j}m_{j}+(1+d_{j})/2)\alpha_{j}\right\}$
and a similar statement holds for $C^{-}(\underline{a},\underline{\alpha},v,x)$ so that $G_{r,\underline{a}}^{\underline{\alpha}}(v)(w,x)^{(-1)^{D}}$ is equal to:
$\prod_{\delta^{\prime}\in C^{+}(\underline{a},\underline{\alpha},v,x)/\mathbb{Z}\gamma}\left(1-e^{2i\pi\left(\frac{x(\delta^{\prime})-w}{x(\gamma)}\right)}\right)\prod_{\delta^{\prime}\in C^{-}(\underline{a},\underline{\alpha},v,x)/\mathbb{Z}\gamma}\left(1-e^{-2i\pi\left(\frac{x(\delta^{\prime})-w}{x(\gamma)}\right)}\right)^{(-1)^{r}}$
@@
$\prod_{Nv^{\prime}\equiv v\mathrm{\penalty 10000\ mod\penalty 10000\ }L}G_{r,\underline{a}}(v^{\prime})(w,x)=G_{r,\underline{a}}(v)(Nw,x)$ (16)
@@ 3.3 The cocycle $\phi_{n,a}$ for unit groups
We now give several examples of such groups $U$. A simple example is given by the subgroup of $\mathrm{SL}_{n}(\mathbb{Z})$ consisting of matrices which stabilise the set $\{\alpha\in L\penalty 10000\ |\penalty 10000\ a(\alpha)\geq 0\}$. This corresponds to the case where linear forms are on the same side of some hyperplane through the origin in [[FHRZ08], Lemma 3.9]. It is clear that such a group satisfies condition (26) and that it is isomorphic to the subgroup:
@@ 4.1 Shintani’s method
From this point forward, we may take two slightly different approaches to the computation of the full partial zeta functions at $s=0$, one of them following closely Shintani’s original strategy [Shi76] and the other following the “signed fundamental domain” strategy in [DyDF14].
@@ 4.2 Shintani’s fundamental domain
In particular, we show (see [Mor26c]) that $\mathcal{K}(V)=\mathcal{L}(V)+\mathcal{K}^{n}(V)$ where $\mathcal{K}^{n}(V)$ the subspace of $\mathcal{K}(V)$ spanned by the functions $c^{\vee}(a_{1},\dots,a_{n})$ for linearly independent $a_{1},\dots,a_{n}$ in $\Lambda$. The proof of such a statement in fact produces an algorithm that expresses any indicator function $c^{\vee}(a_{1},\dots,a_{m})$ modulo $\mathcal{L}(V)$ in terms of functions $c^{\vee}(a_{j,1},\dots,a_{j,n})$ where each of the $a_{j,l}$’s are equal to some $\pm a_{k}$. This gives the following:
@@
where we have factored in the $\mathrm{gcd}$’s of the coefficients. We carried out the computations using the computer software Pari/GP [The23] and found:
@@
Proposition 4.2 is already a great way to express the partial zeta values in terms of these Bernoulli polynomials $B_{n,a_{1},\dots,a_{n}}$ related to the multiple elliptic Gamma functions. However, the explicit computation of this decomposition is quite tedious in general. In addition, the linear forms involved are not quite of the form we hoped for following the discussion at the end of section 3.3. Indeed, we would like to evaluate the Bernoulli polynomials on a cycle whose shape would resemble the cycle described in [[CDG15], section 2.6] which corresponds to the sign fundamental domain decomposition in [DyDF14]. This is the second approach which we carry out in the next section.

### v2

@@ Abstract:
This is the first paper in a series where we study arithmetic applications of the multiple elliptic Gamma functions originated from mathematical physics. The main purpose of this paper is the introduction of a framework for applications of these functions to Hilbert’s 12th problem for general number fields with exactly one complex place following recent work by Bergeron, Charollois and García. Namely, we define geometric families of the multiple elliptic Gamma functions, upgrading the construction carried out by Felder, Henriques, Rossi and Zhu for rank $3$ lattices to lattices of higher ranks. These functions enjoy transformation properties under an action of the special linear group $\mathrm{SL}_{n}(\mathbb{Z})$ for $n\geq 2$ involving some Bernoulli rational functions as their so-called modularity defect. A second purpose of this paper is to use this collection of Bernoulli rational functions to construct (n1)(n-1)-cocycles for specific subgroups of $\mathrm{SL}_{n}(\mathbb{Z})$ associated to groups of units in totally real number fields and use these cocycles to compute partial zeta values at $s=0$.
@@ 1 Introduction
where as usual the notation $\widehat{\tau_{j}}$ indicates that the variable $\tau_{j}$ should be omited. In [FHR+08] Felder, Henriques, Rossi and Zhu enriched the theory of the elliptic Gamma function by introducing geometric families of elliptic Gamma functions associated to arbitrary rank $3$ lattices, offering a comprehensive perspective of the underlying geometric phenomena. Namely, if $L$ is a rank $3$ lattice and $a,b\in\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$ are two primitive linear forms on $L$, they define general functions $\Gamma_{a,b}:\mathbb{C}\times\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{C})\to\mathbb{C}$ using Ruijsenaars’ elliptic $\Gamma$ function as a building block and show that the collection of functions obtained when varying $a,b$ enjoys nice properties under an action of $\mathrm{SL}_{3}(\mathbb{Z})$ on $L$. The two most important properties in that regard are the so-called modular and equivariance properties. For any linearly independent $a,b,c\in\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$, there is a rational function $P_{a,b,c}\in\mathbb{Q}[w](x)$ such that the equality:
@@
In the first part of this article, we upgrade the construction carried out in [FHR+08] to higher degree $G_{r}$ functions (see section 2). For a lattice $\Lambda=\mathrm{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$ of rank $n=r+2\geq 2$ we introduce a collection of geometric functions $G_{n-2,a_{1},\dots,a_{n-1}}$ attached to families of n1n-1 linear forms $a_{1},\dots,a_{n-1}\in\Lambda$ (see Proposition 2.2) which are built using Nishizawa’s $G_{n-2}$ functions. For rank $n=2,3$ lattices, we recover the functions $\theta_{a}=G_{0,a}(0)$ and $\Gamma_{a,b}=G_{1,a,b}(0)$ respectively. In section 2 we show that the collection of $G_{n-2,a_{1},\dots,a_{n-1}}$ functions obtained when varying the linear forms $a_{1},\dots,a_{n-1}$ satisfy modular and equivariance properties similar to those satisfied by the collections of $\theta_{a}$ and $\Gamma_{a,b}$ functions. Namely, our first main result is a general version of formulae (1) and (2) for the $G_{n-2,a_{1},\dots,a_{n-1}}$ functions involving a collection of higher degree Bernoulli rational functions $B_{n,a_{1},\dots,a_{n}}$ (see Definition 2.2).
@@
In applications the linear forms $a_{j}$ will be taken of the form $g_{j}\cdot a$ for some base point $a\in V^{\vee}$ and the invertible matrix $g_{j}$ in some subgroup $U$ of $\mathrm{SL}_{n}(\mathbb{Z})$. At the end of section 3 we give examples of specific subgroups $U$ of $\mathrm{SL}_{n}(\mathbb{Z})$ for which any family $g_{0},\dots,g_{n}\in U$ is such that the family $g_{0}\cdot a,\dots,g_{n}\cdot a$ satisfies the hypothesis of Corollary 1 for any base point $a\in V^{\vee}$. The collection of functions $(\phi_{n,a})_{a}$ thus reduces to a collection of homogeneous (n1)(n-1)-cocycles for $U$. In the arithmetic applications we have in mind to the computation of partial zeta values in totally real number fields at $s=0$, the group $U$ will arise from the group of totally positive units of a given totally real number field. Following [SHI76], [COL88], [DF14] and [CDG15] we will prove that the partial zeta values in totally real number fields may be expressed as combinations of values of the geometric Bernoulli functions $B_{n,a_{1},\dots,a_{n}}$ which appear in the study of the collection of $G_{n-2,a_{1},\dots,a_{n-1}}$ functions:
@@ 2 Geometric families of multiple elliptic Gamma functions
In this section, we recall the properties of the multiple elliptic Gamma functions (the $G_{r}$ functions) and construct geometric families $G_{r,a_{1},\dots,a_{r+1}}$ of these functions attached to a family of $r+1$ linear forms $a_{1},\dots,a_{r+1}$, upgrading the $G_{r}$ functions to collections of equivariant functions for $\mathrm{SL}_{r+2}(\mathbb{Z})$ by adapting the construction of the $\Gamma_{a,b}$ functions in [FHR+08] to higher degrees.
@@ 2.1.3 Distribution relations
$\mathbb{Z}_{>0}-\{0,1/N,\dots,(N-1)/N\}=\mathbb{Z}_{>0}/N,\penalty 10000\ \penalty 10000\ \mathbb{Z}_{\geq 0}+\{0,1/N,\dots,(N-1)/N\}=\mathbb{Z}_{\geq 0}/N.$
@@ 2.2 Geometric $G_{r,\underline{a}}$ functions
In this section we upgrade the construction carried out by Felder, Henriques, Rossi and Zhu to higher degree and define geometric families $G_{r,a_{1},\dots,a_{r+1}}$ parametrised by $r+1$ linear forms of the $G_{r}$ functions that encompass the geometric families $\Gamma_{a,b}$ of the elliptic Gamma function. We then prove Theorem 1, i.e. the modular property for the $G_{r,a_{1},\dots,a_{r+1}}$ functions and their equivariance property under the action of $\mathrm{SL}_{r+2}(\mathbb{Z})$. The construction of our $G_{r,a_{1},\dots,a_{r+1}}$ functions is adapted from the construction of the elliptic $\Gamma_{a,b}$ functions as we use a generalised version of the alternative definition given by [[FHR+08], Proposition 3.5] and then revert the computations to prove that the definition is indeed valid.
@@
We may now define the geometric variants of the $G_{r}$ functions using lemma 2.2 and adapting [[FHR+08], Proposition 3.5] to higher rank lattices:
@@
$G_{r,\underline{a}}^{\underline{\alpha}}(v)(w,x)^{(-1)^{D}}=\prod_{\delta\in F(\underline{a},\underline{\alpha},v)/\mathbb{Z}\gamma}\prod_{\underline{m}\,\geq 0}\Big[\left(1-e^{2i\pi\left(\sum_{j=1}^{r+1}\frac{(d_{j}m_{j}+(1+d_{j})/2)x(\alpha_{j})}{x(\gamma)}-\frac{w+x(\delta)}{x(\gamma)}\right)}\right)\\ \times\left(1-e^{2i\pi\left(\sum_{j=1}^{r+1}\frac{(d_{j}m_{j}+(d_{j}-1)/2)x(\alpha_{j})}{x(\gamma)}+\frac{w+x(\delta)}{x(\gamma)}\right)}\right)^{(-1)^{r}}\Big]$
Let us denote by $C^{+}(\underline{a},\underline{\alpha},v,x)$ the set of $\delta^{\prime}\in v+L$ satisfying for all $1\leq j\leq r+1$:
@@
and similarly denote by $C^{-}(\underline{a},\underline{\alpha},v,x)$ the set of $\delta^{\prime}\in v+L$ satisfying for all $1\leq j\leq r+1$:
@@
Consider $\delta^{\prime}\in C^{+}(\underline{a},\underline{\alpha},v,x)$. If $d_{j}=1$ then $a_{j}(\delta^{\prime})\geq 0$ so that performing Euclidian division by $a_{j}(\alpha_{j})$ gives a unique integer $m_{j}\geq 0$ satisfying $0\leq a_{j}(\delta^{\prime}-m_{j}\alpha_{j})<a_{j}(\alpha_{j})$. On the contrary, if $d_{j}=-1$ then $a_{j}(\delta^{\prime})<0$ and there exists a unique integer $m_{j}>0$ such that $0\leq a_{j}(\delta^{\prime}+m_{j}\alpha_{j})<a_{j}(\alpha_{j})$. Then for all $1\leq k\leq r+1$:
$0\leq a_{k}\left(\delta^{\prime}-\sum_{j=1}^{r+1}d_{j}m_{j}\alpha_{j}\right)<a_{k}(\alpha_{k})$
and thus $\delta^{\prime}-\sum_{j=1}^{r+1}d_{j}m_{j}\alpha_{j}\in F(\underline{a},\underline{\alpha},v)$. This shows that the cone $C^{+}(\underline{a},\underline{\alpha},v,x)$ can be written as a disjoint union:
$C^{+}(\underline{a},\underline{\alpha},v,x)=\bigcup_{\delta\in F(\underline{a},\underline{\alpha},v)}\bigcup_{\underline{m}\,\geq 0}\left\{\delta+\sum_{j=1}^{r+1}(d_{j}m_{j}+(d_{j}-1)/2)\alpha_{j}\right\}.$
A similar argument applied to $C^{-}(\underline{a},\underline{\alpha},v,x)$ gives the decomposition:
$C^{-}(\underline{a},\underline{\alpha},v,x)=\bigcup_{\delta\in F(\underline{a},\underline{\alpha},v)}\bigcup_{\underline{m}\,\geq 0}\left\{\delta-\sum_{j=1}^{r+1}(d_{j}m_{j}+(d_{j}+1)/2)\alpha_{j}\right\}.$
Thus, the expression $G_{r,\underline{a}}^{\underline{\alpha}}(v)(w,x)^{(-1)^{D}}$ is equal to:
$\prod_{\delta^{\prime}\in C^{-}(\underline{a},\underline{\alpha},v,x)/\mathbb{Z}\gamma}\left(1-e^{-2i\pi\left(\frac{w+x(\delta^{\prime})}{x(\gamma)}\right)}\right)\prod_{\delta^{\prime}\in C^{+}(\underline{a},\underline{\alpha},v,x)/\mathbb{Z}\gamma}\left(1-e^{2i\pi\left(\frac{w+x(\delta^{\prime})}{x(\gamma)}\right)}\right)^{(-1)^{r}}$
@@
$\prod_{Nv^{\prime}\equiv v\mathrm{\penalty 10000\ mod\penalty 10000\ }L}G_{r,\underline{a}}(v^{\prime})(w,x)=G_{r,\underline{a}}(v)(Nw,x).$ (16)
@@ 3.3 The cocycle $\phi_{n,a}$ for unit groups
We now give several examples of such groups $U$. A simple example is given by the subgroup of $\mathrm{SL}_{n}(\mathbb{Z})$ consisting of matrices which stabilise the set $\{\alpha\in L\penalty 10000\ |\penalty 10000\ a(\alpha)\geq 0\}$. This corresponds to the case where linear forms are on the same side of some hyperplane through the origin in [[FHR+08], Lemma 3.9]. It is clear that such a group satisfies condition (26) and that it is isomorphic to the subgroup:
@@ 4.1 Shintani’s method
From this point forward, we may take two slightly different approaches to the computation of the full partial zeta functions at $s=0$, one of them following closely Shintani’s original strategy [SHI76] and the other following the “signed fundamental domain” strategy in [DF14].
@@ 4.2 Shintani’s fundamental domain
In particular, we show (see [MORpe]) that $\mathcal{K}(V)=\mathcal{L}(V)+\mathcal{K}^{n}(V)$ where $\mathcal{K}^{n}(V)$ the subspace of $\mathcal{K}(V)$ spanned by the functions $c^{\vee}(a_{1},\dots,a_{n})$ for linearly independent $a_{1},\dots,a_{n}$ in $\Lambda$. The proof of such a statement in fact produces an algorithm that expresses any indicator function $c^{\vee}(a_{1},\dots,a_{m})$ modulo $\mathcal{L}(V)$ in terms of functions $c^{\vee}(a_{j,1},\dots,a_{j,n})$ where each of the $a_{j,l}$’s are equal to some $\pm a_{k}$. This gives the following:
@@
where we have factored in the $\mathrm{gcd}$’s of the coefficients. We carried out the computations using the computer software Pari/GP [18] and found:
@@
Proposition 4.2 is already a great way to express the partial zeta values in terms of these Bernoulli polynomials $B_{n,a_{1},\dots,a_{n}}$ related to the multiple elliptic Gamma functions. However, the explicit computation of this decomposition is quite tedious in general. In addition, the linear forms involved are not quite of the form we hoped for following the discussion at the end of section 3.3. Indeed, we would like to evaluate the Bernoulli polynomials on a cycle whose shape would resemble the cycle described in [[CDG15], section 2.6] which corresponds to the sign fundamental domain decomposition in [DF14]. This is the second approach which we carry out in the next section.